### v1

@@ Abstract
We extend a previous result [Sutter et al, J. Phys. A: Math. Theor. 57, 475202 (2024)] to give an explicit form of the set of $v$-representable densities on the one-dimensional torus with any fixed number of particles in contact with a heat bath at finite temperature. The particle interaction has to satisfy some mild assumptions but is kept entirely general otherwise. For densities, we consider the Sobolev space $H^{1}$ and exploit the convexity of the functionals. This leads to a broader set of potentials than the usual $L^{p}$ spaces and encompasses distributions. By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is guaranteed. This yields $v$-representability and it is demonstrated that the given set of $v$-representable densities is even maximal.
@@ I Introduction
Density-functional theory (DFT) is based on the one-to-one relation between densities and potentials. The famous work of Hohenberg and Kohn [1] introduced an energy functional in terms of the density which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’. The lack of knowledge about this property can be circumvented by using the constrained-search functional as introduced by Percus [2] and Levy [3]. This functional is defined for all densities coming from a $N$-particle wave function ($N$-representability). However, for the method of Kohn and Sham [4] it is still crucial to solve the $v$-representability problem. The idea of this method is to reproduce the ground-state density of an interacting Hamiltonian by a non-interacting Hamiltonian with another external potential. Thus, the density needs to be $v$-representable for an interacting system but at the same time also for a non-interacting one. A possible extension is to allow for mixed states. Levy [5] and Lieb [6] showed that there are ensemble ground-state density which can not be realized by a pure ground-state. Hence the set of $v$-representable densities is extended by allowing for ensemble states and not only pure states. In [7], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$. Shortly after, Carvalho Corso [8, 9] showed that for an odd number of particles on the one-dimensional torus each ground-state density of a Hamiltonian with a potential in $H^{-1}(\mathbb{T})$ is non-zero. This result completes the full characterization of $v$-representable densities on the one-dimensional torus. Apart from that it was shown that under the same condition for the number of particles, the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities. Additionally, Carvalho Corso [10] presented these results for non-interacting systems but without any requirement on the number of particles.
@@
Already in 1965 Mermin [11] extended the work of Hohenberg and Kohn [1] to non-zero temperatures. However the same question about $v$-representability appears. Therefore we are interested in extending our recent work [7] from zero temperature to elevated temperatures. First results about this problem were presented by Chayes et al. [12] for infinite lattice systems. Eschrig [13] formally argued that for the grand canonical ensemble on a $3$-dimensional torus the thermal universal functional is differentiable at $v$-representable densities and that these densities are strictly positive. However, he does not give an explicit expression for the set of $v$-representable densities. In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the universal functional (the set of physical densities) is not open in $L^{3}$. In particular, arbitrary close to any physical density there is n density which takes negative values. Hence the universal functional is discontinuous at each physical density. This problem is already extensively studied by Lammert [14] for the zero temperature case. While Lammert is able to show differentiability at certain densities, he does not characterize the set of $v$-representable densities, a problem finally solved on the one-dimensional torus in Sutter et al. [7]. In our previous and the present work, the problems with openness and continuity are circumvented by considering the $H^{1}$ topology instead of the $L^{3}$ topology. Then, the set of $v$-representable densities is open and the universal functional is continuous on that set. Hence, we can investigate Gâteaux differentiability. Another important difference to Eschrig’s work is that we allow for a broader class of potentials and we show that this is even necessary for $v$-representability.
@@
In this work we consider the same class of systems as in Sutter et al. [7] but at elevated temperatures. This means that there are a fixed number of particles on the one-dimensional torus and that we allow for a whole class of interactions including $W=0$. Elevated temperature means that besides the Hamiltonian an additional entropic term has to be considered in the variational principle. In particular, this means that instead of the usual internal energy functional we work with the Helmholtz functional. The minimizer of the Helmholtz functional is explicitly known in terms of the Hamiltonian and it is called Gibbs state. As in [15; 16] the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the universal functional for all interactions in the considered class. Additionally, we show that each density coming from a Gibbs state is strictly positive as Carvalho Corso [8, 9] showed for zero temperature and a odd number of particles. Hence we get the full characterization of $v$-representable densities. The reason for these additional results compared to the zero temperature case with even number of particles, is that the minimizer of Helmholtz functional occupies not only the ground state of the Hamiltonian but also all excited states. Degenerate states get equal weights which yields a unique minimizer of the Helmholtz functional and strict convexity of the Helmholtz free energy functional (in terms of potentials). We formulate the main result of this work.
@@
Define the following set of densities on the one-dimensional torus $\mathbb{T}$ and the space of (distributional) one-body potentials,
$\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (1a)
$\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (1b)
Then for every $\rho\in\mathscr{X}_{>0}$ there is an equivalence class $v]\in\mathscr{X}^{*}$ with corresponding potential $V=\sum_{j=1}^{N}v(x_{j})$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V$, with $W$ any (interaction) operator that fulfills the KLMN condition (see Eq. (3)). On the other hand, each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. Moreover, the thermal universal functional $F^{\beta}_{\mathrm{DM}}(\rho)$ (defined in Eq. (15b)) is differentiable if and only if $\rho\in\mathscr{X}_{>0}$. This implies that every $\rho\in\mathscr{X}_{>0}$ is even uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
The ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [1]. To prove the stated theorem above, we start by introducing the basic setting for a $N$-particle system on a one-dimensional torus and the relevant spaces for the densities and the potentials (Section II). Afterwards, we introduce the Helmholtz functional. We present known result about this functional and about the Gibbs state together with the introduction of the thermal universal functional (Section III). Then we show the relation between $v$-representability and subgradients (Section IV) followed by an application of the preceding results to conclude $v$-representability (Section V). We then prove the converse statement, that each Gibbs state has indeed a density in the $v$-representable set, to get a full characterization of the set of $v$-representable densities (Section VI). Lastly, we show that the universal functional is Gâteaux differentiable (Section VII). We conclude with the proof of Theorem 1 and an outlook (Section VIII).
@@ II One-dimensional torus setting
We consider a system at an elevated temperature $T$ consisting of $N$ spin $\frac{1}{2}$-particles. The particles are on the one-dimensional torus $\mathbb{T}$, meaning that they are confined to the interval $[0,1]$ where the points 0 and $1$ are identified with each other. The underlying Hilbert space is then
$\mathcal{H}=\bigl{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}.$ (2)
As noted, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$ where $H_{0}$ consists of the kinetic energy operator and an interaction term and where $V$ is the potential. We can generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates. The interaction is kept general but positive and it needs to fulfill the KLMN condition (Reed and Simon [17], Th. X.17) with respect to the kinetic energy operator, i.e. for each interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that the following holds for each element in the form domain $Q(-\frac{1}{2}\Delta)$
@@
This condition guarantees that the Hamiltonian $H_{0}=-\tfrac{1}{2}\Delta+W$ remains self-adjoint with the same form domain $Q(H_{0})=Q(-\frac{1}{2}\Delta)$ if we add an interaction to the kinetic energy operator (which is self-adjoint). In Sutter et al. [7, Appendix] this condition was shown for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$. Here and in the following we use the notion of quadratic forms [18, Sec. VIII.6] instead of linear operators. For the kinetic energy operator the form domain is given by $Q(-\frac{1}{2}\Delta)=\bigl{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}$ with the Sobolev space $H^{1}(\mathbb{T})=\{\Psi\in L^{2}(\mathbb{T})\mid\nabla\Psi\in L^{2}(\mathbb{T})\}$ (like Eq. 1a this means a weak derivative). We get by integration by parts
$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle.$ (4)
$\mathscr{W}=\{\Psi\in(H^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (5)
The one-particle density $\rho_{\Psi}$ is obtained by integrating out all but one of the spatial coordinates of a wave function squared as well as summing over all spin components, i.e.
In the following theorem we collect results from [6] and [7].
@@
We define the following set of densities
$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\}.$ (7)
@@
    There exist two constants $C_{1},C_{2}>0$ such that for all $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and the following estimate holds
@@
    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}$ and where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
Lieb [6] showed that the set $\mathscr{I}$ is convex. This implies that (finite) convex combinations of elements in $\mathscr{I}$ lie again in $\mathscr{I}$. Motivated by the observations for the zero temperature case [7] we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of $\mathscr{X}$. This space contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if they differ only by a constant, i.e., like in Eq. (1b),
@@
The dual pairing of an element $v=f+\nabla g\in H^{-1}(\mathbb{T})$ with an element $\varphi\in H^{1}(\mathbb{T})$ is given by $v(\varphi)=\langle v,\varphi\rangle=\langle f,\varphi\rangle-\langle g,\nabla\varphi\rangle$. Using the dual space of $\mathscr{X}$ for the external potentials guarantees that the potential energy is finite. That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator is shown in [7] by using the KLNM theorem (Reed and Simon [17], Th. X.17).

### v2

@@ Abstract
We extend a previous result [Sutter et al, J. Phys. A: Math. Theor. 57, 475202 (2024)] to give an explicit form of the set of $v$-representable densities on the one-dimensional torus with any fixed number of particles in contact with a heat bath at finite temperature. The particle interaction has to satisfy some mild assumptions but is kept entirely general otherwise. Since we find the necessary domain for densities to be the Sobolev space $H^{1}$, this leads to a broader dual space for potentials than the usual $L^{p}$ spaces and encompasses distributions. By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is achieved. This is equivalent to unique $v$-representability and we can prove that the given set of $v$-representable densities is even maximal.
@@ I Introduction
Density-functional theory (DFT) is based on the one-to-one relation between densities and potentials. The famous work of Hohenberg and Kohn [21] introduced an energy functional in terms of the density, which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’. The lack of knowledge about this property, termed the “$v$-representability problem”, can be circumvented by using the constrained-search functional as introduced by Percus [33] and Levy [26]. This functional is defined for all densities coming from an $N$-particle wave function with finite kinetic energy ($N$-representability). However, for the method of Kohn and Sham [23] it is still crucial to solve the $v$-representability problem. The idea of this method is to reproduce the ground-state density of an interacting Hamiltonian by a non-interacting Hamiltonian with another external potential. Thus, the density needs to be simultaneously $v$-representable for an interacting and a non-interacting system. A possible extension is to allow for mixed states. Levy [27] and Lieb [28] showed that there are ensemble ground-state densities which cannot be realized by a pure ground state. Hence the set of $v$-representable densities is extended by allowing for ensemble states and not only pure states. In Sutter et al. [38], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$. Shortly after, Carvalho Corso [7, 8] showed that for an odd number of particles on the one-dimensional torus each ground-state density of a Hamiltonian with a potential in $H^{-1}(\mathbb{T})$ is non-zero. This result completes the full characterization of $v$-representable densities in this case. The same references also show that the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities. Additionally, Carvalho Corso [9] presents these results for non-interacting systems but without any requirement on the number of particles.
@@
Already in 1965 Mermin [30] extended the work of Hohenberg and Kohn [21] to non-zero temperatures. Recently, there has been renewed, broad interest in the field of thermal DFT, ranging from the application of DFT to warm dense matter [22; 36; 18] and Hubbard models [31] to more theoretical work concerning exact conditions [5], the adiabatic connection [19], a rigorous proof of the Hohenberg–Kohn theorem for systems with temperature [16], and the formulation of Entropic DFT [39]. For this reason, we want to try and answer the same question about $v$-representability for the case of thermal ensembles and hereby extend our recent work [38] from zero temperature to elevated temperatures. First results about this problem were achieved by Chayes et al. [10] for infinite lattice systems. Eschrig [13] formally argued that for the grand canonical ensemble on a $3$-dimensional torus the thermal universal functional is differentiable at $v$-representable densities and that these densities are strictly positive. However, he does not give an explicit expression for the set of $v$-representable densities. In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the thermal universal functional (the set of physical densities) is not open in $L^{3}$. The same problem persists for the usual universal functional of DFT. In particular, arbitrary close to any physical density there is a density that takes negative values. Hence the universal functional is discontinuous at each physical density. This problem was previously studied extensively by Lammert [25] for the zero temperature case. While Lammert was able to show differentiability at certain densities by considering a Sobolev space topology, he did not characterize the set of all $v$-representable densities, a problem finally solved in Sutter et al. [38] on the one-dimensional torus. In our previous and the present work, the problems with openness and continuity are circumvented by considering the finer $H^{1}$ topology instead of the $L^{3}$ topology. A topology is finer than another if it has more open sets, or equivalently, less convergent sequences. In one dimension it holds in terms of norms that for any $2\leq p\leq\infty$ there is a $C>0$ such that for all $f\in H^{1}$ one has
$\|f\|_{L^{p}}\leq C\|f\|_{H^{1}}$ (1)
and one says that $H^{1}$ is continuously embedded into $L^{p}$ [1, Th. 4.12, Part I.A]. Consequently, in one dimension the $H^{1}$ topology is finer than all $L^{p}$ topologies, $p\geq 2$. This leads to the set of strictly positive densities in $H^{1}$ being open and to a universal functional that is continuous on that set. Hence, we can subsequently investigate Gâteaux differentiability, the property that for a functional $f:X\to\mathbb{R}$ on a Banach space $X$ at every point $x\in U$ from an open domain $U\subseteq X$ the functional derivative
$\nabla f(x)=\lim_{\tau\to 0}\frac{f(x+\tau y)-f(x)}{\tau}$ (2)
exists for all directions $y\in X$.
@@
In this work, we study the same class of systems as in Sutter et al. [38] but at elevated temperatures. This means that we consider a fixed number of fermionic particles on the one-dimensional torus and a very general class of interactions, including of course $W=0$. Elevated temperature means that besides the Hamiltonian an additional entropic term has to be added in the variational principle, and that instead of the usual energy we work with the Helmholtz free energy. The minimizer of the Helmholtz free energy is explicitly known in terms of the Hamiltonian and it is called Gibbs state. As demonstrated before [37; 17], the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the thermal universal functional for all interactions in the considered class. Additionally, we show that each density coming from a Gibbs state is strictly positive. The same is true for zero temperature and an odd number of particles [7; 8]. Hence we get the full characterization of $v$-representable densities. The reason for these additional results is that the minimizer of the Helmholtz free energy does not only occupy the ground state of the Hamiltonian but partially also all excited states. Degenerate states get equal weights and this yields a unique minimizer and strict convexity of the Helmholtz functional (in terms of potentials). We now formulate the main result of this work.
@@
Define a set of densities and a space of (distributional) one-body potentials on the one-dimensional torus $\mathbb{T}$, $\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (3a) $\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (3b) Then for every $\rho\in\mathscr{X}_{>0}$ there is a unique potential $v$ with equivalence class $v]\in\mathscr{X}^{*}$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V,\quad V=\sum_{j=1}^{N}v(x_{j}),$ (4) where $W$ is any (interaction) operator that fulfills the KLMN condition (see Eq. (10)). On the other hand, each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. Moreover, the thermal universal functional $F^{\beta}_{\mathrm{DM}}(\rho)$ (defined in Eq. (22b)) is Gâteaux differentiable if and only if $\rho\in\mathscr{X}_{>0}$. This implies that for a fixed interaction $W$, every $\rho\in\mathscr{X}_{>0}$ is uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
The property of ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [21; 15]. To prove the stated theorem above, we start by introducing the basic setting for an $N$-particle system on a one-dimensional torus and the relevant spaces for densities and potentials (Section II). Afterwards, we introduce the Helmholtz free energy and present known results about it and its unique minimizer, the Gibbs state, together with the definition of the thermal universal functional (Section III). Next, we show the non-emptiness of the subdifferential of the thermal universal functional that is a necessary condition for $v$-representability (Section IV). This is followed by an application of the preceding results to conclude $v$-representability (Section V). We then prove the converse statement, that each Gibbs state indeed yields a density in $\mathscr{X}_{>0}$, to achieve a full characterization of the set of $v$-representable densities (Section VI). Lastly, we show that the thermal universal functional is Gâteaux differentiable (Section VII). We conclude with the proof of Theorem 1 that is just a collection of previous results and an outlook (Section VIII).
Table 1: Notation Table
Symbol Description
$\mathbb{T}$ one-dimensional torus
$L^{p}$ real Lebesgue spaces
$L^{p}_{\mathbb{C}}$ complex Lebesgue spaces
$H^{1}$ real Sobolev space $W^{1,2}$
$H^{1}_{\mathbb{C}}$ complex Sobolev space $W^{1,2}$, Eq. (6)
$H^{-1}$ topological dual of the real Sobolev space
$\mathcal{H}$ antisymmetric Hilbert space for $N$ particles, Eq. (5)
$\mathscr{W}$ antisymmetric wave function space based on $H^{1}_{\mathbb{C}}$, Eq. (8)
$\mathscr{D}$ set of density matrices, Eq. (19)
$\mathscr{I}$ set of physical densities, Eq. (12)
$\mathscr{X}$ affine space of normalized densities, Eq. (13)
$\mathscr{X}_{>0}$ strictly positive subset of the affine space, Eq. (3a)
$\mathscr{X}^{*}$ topological dual of the affine space, Eq. (3b)
$\Omega^{\beta}_{v}$ Helmholtz free energy of a density matrices, Eq. (17)
$\Omega^{\beta}$ Helmholtz functional on potentials, Eq. (18)
$F^{\beta}_{\mathrm{DM}}$ thermal universal functional, Eq. (22b)
@@ II One-dimensional torus setting
We consider a system at an elevated temperature $T$ consisting of $N$ spin $\frac{1}{2}$-particles. Choosing a different spin as an internal degree of freedom (or none at all) will not change any of the following results, we just use spin $\frac{1}{2}$ and antisymmetric wave functions to relate more closely to the usual systems considered in quantum chemistry. The particles are on the one-dimensional torus $\mathbb{T}$, meaning that they are confined to the interval $[0,1]$ where the points 0 and $1$ are identified with each other. The underlying complex Hilbert space is then
$\mathcal{H}=\bigl(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}.$ (5)
We also define the complex Sobolev space on the torus
$H_{\mathbb{C}}^{1}(\mathbb{T})=\{f\in L^{2}_{\mathbb{C}}(\mathbb{T})\mid\nabla f\in L^{2}_{\mathbb{C}}(\mathbb{T})\}$ (6)
(like in Eq. (3a) this means a weak derivative) and the form domain of the kinetic energy operator
$Q(-\tfrac{1}{2}\Delta)=\bigl(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}.$ (7)
$\mathscr{W}=\{\Psi\in(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (8)
The one-particle density $\rho_{\Psi}$ for any $\Psi\in\mathscr{W}$ is obtained by integrating out all but one of the spatial coordinates from the modulus squared of a wave function summed over all spin components,
As already stated in Theorem 1, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$, where $H_{0}$ consists of the kinetic energy operator and an interaction term. We generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates. The interaction is kept general and just needs to fulfill the KLMN condition [34, Th. X.17] with respect to the kinetic energy operator. This means that for a fixed interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that for each $\Psi\in Q(-\frac{1}{2}\Delta)$ it holds
@@
This condition guarantees that the Hamiltonian $H_{0}=-\frac{1}{2}\Delta+W$ is self-adjoint with the same quadratic form domain as the kinetic energy operator. In Sutter et al. [38, Appendix] this condition was shown to hold for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$. Here and in the following we use the notion of quadratic forms instead of linear operators [35, Sec. VIII.6]. For the kinetic energy operator this is achieved directly by integration by parts
$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle=T(\Psi).$ (11)
In the following theorem we collect some important previous results [28; 38].
@@
We define the following set of physical densities,
$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\},$ (12)
and an affine space of normalized densities in the real Sobolev space $H^{1}(\mathbb{T})$,
$\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}.$ (13)
@@
    There exist two constants $C_{1},C_{2}>0$ such that for every $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and fulfills the following estimate,
@@
    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
Lieb [28] showed that the set $\mathscr{I}$ is convex. This implies that (finite) convex combinations of elements in $\mathscr{I}$ lie again in $\mathscr{I}$. Motivated by the observations for the zero temperature case [38], we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of the subspace $\{\rho\in H^{1}(\mathbb{T})\mid\int\rho=0\}$ that is isomorphic to $\mathscr{X}$. This space $\mathscr{X}^{*}$ contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if and only if they differ only by a constant, i.e., like in Eq. (3b),
@@
The dual pairing of an element $v=f+\nabla g\in H^{-1}(\mathbb{T})$ with an element $\varphi\in H^{1}(\mathbb{T})$ is given by $v(\varphi)=\langle v,\varphi\rangle=\langle f,\varphi\rangle-\langle g,\nabla\varphi\rangle$. Using $\mathscr{X}^{*}$ as potential space automatically guarantees that the potential energy is finite. That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator was shown in Sutter et al. [38, Th. 18] by using the KLMN theorem [34, Th. X.17].